Find all the maxima and minima of the
function f given by
f(x)= integration of [3(t-2)(t-3)^3 + 2(t-2)^2 (t-3)^2] dt. In the limit 2 to x
1
Expert's answer
2019-05-09T13:40:22-0400
"f(x)=\displaystyle\int_{2}^x[3(t-2)(t-3)^3+2(t-2)^2(t-3)^2]dt"

Domain: "\R"


"f'(x)=[3(x-2)(x-3)^3+2(x-2)^2(x-3)^2]\cdot1-0="

"=(x-2)(x-3)^2(3x-9+2x-4)=(x-2)(x-3)^2(5x-13)"

Find the critical number(s):


"f'(x)=0=>(x-2)(x-3)^2(5x-13)=0"

"x_1=2, x_2={13 \over 5}, x_3=3"

First Derivative Test


"If\ x<2, f'(x)>0, f(x)\ increases"




"If\ 2<x<2.6, f'(x)<0, f(x)\ decreases"

"If\ 2.6<x<3, f'(x)>0, f(x)\ increases"

"If\ x>3, f'(x)>0, f(x)\ increases"

"\int[3(t-2)(t-3)^3+2(t-2)^2(t-3)^2]dt="

"=3\int(t-3)^4dt+3\int(t-3)^3dt+"

"+2\int(t-3)^4dt+4\int(t-3)^3dt+2\int(t-3)^2dt="

"=(t-3)^5+{7 \over 4}(t-3)^4+{2 \over 3}(t-3)^3+C"

"f(x)=\displaystyle\int_{2}^x[3(t-2)(t-3)^3+2(t-2)^2(t-3)^2]dt="

"=(x-3)^5+{7 \over 4}(x-3)^4+{2 \over 3}(x-3)^3-{1 \over 12}"

"f(2)=0"

"f(2.6)=-0.09144"

The function f(x) has a local maximum with value of at "x=2."

The function f(x) has a local minimum with value of "(-0.09144)" at "x=2.6."